Thank you guys!
If f(x) = 2x + 3 and g(x) = x2 + 1,find f(g(3)).
A:23
B:47
C:82
D:90

Answer by FrankM(1040)

This problem take 2 steps,
g(x) = x^2 + 1
g(3)= 3^2+1 = 10
f(x) = 2x + 3
f(7) = 2*10 + 3 =23